A Major League Baseball team signs a free agent player to a new contract, which includes a $2 million signing bonus. If the player puts the money into an account earning 9%, compounded quarterly, how much will be in the account at the end of 5 years when his contract expires? Show work in excel
Answer #1

Money in account at the end of 5 years is calculated using FV function in Excel :

rate = 9%/4 (Quarterly rate of return = annual rate of return / 4)

nper = 5 * 4 (total number of quarterly periods in 5 years = 5 * 4)

pmt = 0 (periodic deposit into account is zero)

pv = -2000000 (Amount put into account. This is entered with a negative sign because it is a payment into the account)

FV is calculated to be $3,121,018.40

X foc =FV(9%/4,5*4,0,-2000000) D E B C 1 $3,121,018.40

Money in account at the end of 5 years is $3,121,018.40